OpenStudy (anonymous):

describe and correct the error in finding the difference shown -4 - (-5)=-4+(-(-5)) =14+5 =-1

Directrix (directrix):

- 4 - (-5) = - 4 + (-(-5)) = <<< -4 + (+5) = >>> 14+5 = <<< there is no 14; I think you read - 4 as 14>>> <<< - 4 + 5 = >>> -1 <<<< should be 1 >>> --> -4 + 5 = 1 1 So the work goes like this: - 4 - (-5) = - 4 + (-(-5)) = -4 + (+5) = -4 +5 = +1